OpenStudy (anonymous):

I award medals!! Please give me an example of a function with two real rational solutions. I would like the function to be f(x) please. I am trying to grasp what my math lesson is teaching me but the examples aren't that great.

OpenStudy (anonymous):

@timaashorty @marylou004 @phi

OpenStudy (phi):

fx)= (x-1)(x-2) has two solutions to f(x)=0 but can you post what the lesson is saying ?

OpenStudy (anonymous):

but I don't really think that is what the question is looking for.

OpenStudy (anonymous):

I reread the lesson and cant find anything on it. I looked at the examples but they don't have equations like the one you posted

OpenStudy (phi):

can you post one of the examples?

OpenStudy (phi):

I cannot tell for sure, but it looks like they are translating a quadratic from standard form to "vertex form" see http://www.mathwarehouse.com/geometry/parabola/standard-and-vertex-form.php if so, you would do this: \[ f(x)= x^2+12x +26 \\\text{ change f(x) to y}\\ y= x^2+12x +26 \\ \text{add } \left(\frac{12}{2}\right)^2 \text{ to both sides}\\ y+36= x^2+12x +36 +26 \\ \] rewrite \( x^2+12x +36 \) as \( (x+6)^2 \) \[ y+36= (x+6)^2+26 \] add -36 to both sides to get vertex form \[ y= (x+6)^2 -10 \]

OpenStudy (anonymous):

oh okay. That's easy enough.

OpenStudy (anonymous):

Thank you :-)